HUBBLE SPACE TELESCOPE DAILY REPORT #1670 PERIOD COVERED: 0000Z (UTC) 07/08/96 - 0000Z (UTC) 07/09/96 Daily Status Report as of 191/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6354 (Measuring Luminosity Evolution in Z =3D = 0.3 Field Galaxies) The WF/PC-2 was used to make snapshot observations of the galactic = cluster Gal-Cluster-1 located at a distance of Z =3D 0.3. The targets were drawn from a much larger pool of galaxies for which the = internal kinematics have been measured from an extensive ground based = program. A status buffer error message was received at 190:02:25:47Z = which indicated that a failure of the FOS/RD binary target acquisition = scheduled for 02:20:56Z. No other anomalies were noted. 1.2 Completed FOS/RD 6801 (C and N Production and Pollution Mechanism = in Low-Metallicity) The Faint Object Spectrograph (Red side) was used to make = spectrographic observations of the dwarf galaxy UM469 to measure its N, O = and C emission lines in its H II regions. The observations were completed = as planned and no anomalies were reported. 1.3 Completed Three Sets of WF/PC-2 5971 (HST Medium Deep Survey) The WF/PC-2 was used to make a long exposure medium deep sky = survey observations at latitudes larger than 10 degrees taken in parallel = with the prime science observations. The observations were executed as = scheduled, and no anomalies were noted. 1.4 Completed GHRS/1/2 6594 (White Dwarf Companions of Barium and CH = Peculiar Stars) The High Resolution Spectrograph (Sides One and Two ) were used to = attempt to make spectrographic observations of the white dwarf of the = Barium and CH Peculiar class star HD16458-B. This will test the theory = that abundance anomalies in this star caused by mass transfer from the = precursor of the white dwarf. The observations were executed as = scheduled, and no anomalies were noted. 1.5 Completed COSTAR/U 6895 (FOC Cycle 6 PSF Monitoring) The Faint Object Camera/Costar were used to make observations of = object 01-001 (aka, white dwarf star BPM16274) in order to check out the = Point Spread Function (PSF) and focus of the system. The COSTAR DOB was = moved 3.0 microns and was completed at 190/1354Z. The current position is = -0.65 mm. The observations were executed as scheduled, and no problems = were reported. 1.6 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.7 Completed Two Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: = Dark) The WF/PC-2 was used to make internal dark calibration images. The = observations were completed as scheduled, and no problems were reported. 1.8 Completed FOS/BL 6465 (The UV Upturn of Super-Metal-Rich Globular = Clusters) The Faint Object Spectrograph (Blue Side) was used to make = ultraviolet spectrographic observations of the super-metal-rich globular = cluster NGC-224-G17 which is located within the Andromeda Galaxy (aka, = M-31) in order to observe the spectral energy distribution of this = cluster. The objective is to determine whether the strong ultraviolet = fluxes seen in the galaxy can also be seen in this cluster. The = observations were executed as scheduled, and no problems were reported 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Operations Notes: Normal operations, no problems to report. The COSTAR DOB was moved 3.0 microns and was completed at 190/1354Z. The = current position is -0.65 mm, next move will nominally take place in = October or November of this year and will be synchronized with a = secondary mirror move. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH